### v3

# Targeted Nakamoto: A Bitcoin Protocol to Balance Network Security and Energy Consumption

Abstract
In a Proof-of-Work blockchain such as Bitcoin mining hashrate is increasing in the block reward. An increase in hashrate reduces network vulnerability to attack (a reduction in security cost) while increasing carbon emissions and electricity cost (an increase in externalities cost). This implies a tradeoff in total cost at different levels of hashrate and the existence of a hashrate interval where total cost is minimized. Targeted Nakamoto is a Proof-of-Work protocol augmentation that incentivizes miners to home in on a target hashrate interval. When hashrate is above target a ceiling is placed on the block reward a miner can receive. When hashrate is below target a floor is placed underneath the miner’s block reward. Monetary neutrality is maintained by a proportional increase in spending potential among addresses holding UTXO’s to match a deduction from total block reward when the ceiling is operative and a proportional reduction in spending potential among addresses holding UTXO’s to match an increase over the total block reward when the floor is binding.
Bitcoin faces two, diametrically opposed and temporally divergent, existential risks. Today, the massive energy consumed by Bitcoin mining is a focus of a political backlash aimed at de-legitimizing Bitcoin and restricting its use. In the future, as the minting of new UTXO’s to reward miners decreases toward zero, the revenue from mining will likely shrink, causing miners to exit and energy consumption to contract. The security of the network may be compromised by the low cost of an attack. In effect, Bitcoin’s energy consumption will be too low. Targeted Nakamoto is a modification to the Bitcoin protocol that is designed to avert these twin potential disasters.
footnotetext: I wish to thank Madars Virza for his aid and advice in developing the core ideas of this paper; Neha Narula for helping to refine and clarify the presentation; thoughtful comments from participants at the Chaincode Labs research seminar and Mark Serrahn for producing the figures herein. Kristain Praizner wrote the code for the hashrate control algorithm and implemented it on an API that is a companion to this paper.
Keywords— Bitcoin, Blockchain, Mechanism Design, Network Security
1 Introduction
Proof-of -Work (PoW) blockchain cryptocurrencies like Bitcoin require the application of computing power by miners to operate the network. Miners assemble blocks and compete to solve a puzzle set by the code. The number of puzzle guesses (one hash per guess) a mining computer makes in a specified interval of time is its hashrate, which consumes electricity. Miners pay the direct cost of electricity. There are, in addition, two externalities created by mining activity. One externality is the vulnerability of the blockchain network to attack. A measure of vulnerability is the cost of forking the blockchain, which is related to the hashrate of active miners. The higher is hashrate, the higher is the cost of forking a PoW blockchain, which implies that network security is increasing in hashrate (equivalently, security cost is decreasing in hashrate). The other externality arises from the electricity consumed by hashrate. Miner demand for electricity increases both the price and aggregate consumption of electricity. The former is a cost imposed on competing users of electricity. The latter generates carbon emissions, which has a social cost. These electricity related costs are increasing in hashrate. The interaction of security risk and electricity consumption implies a tradeoff in total cost, defined as the sum of the two externalities costs, at different levels of hashrate. An increase in hashrate reduces the vulnerability of the network (equivalently, security cost is decreased), while it increases carbon emissions and creates upward pressure on the price of electricity. There may be a point or interval of hashrate in which total cost is minimized. Current PoW protocols, which I refer to as ”Nakamoto”, do not restrict hashrate. An increase in the value of the block reward creates an incentive for miners to apply more hashrate. A decrease in the value of the block reward has the opposite effect. Equilibrium hashrate can range from zero to unbounded. I focus on Bitcoin in this paper, although the protocol applies to any PoW cryptocurrency. Figure 4 shows the tight correlation over time between percentage change in the USD value of the Bitcoin block reward and the percentage change in energy consumption from Bitcoin mining.
Figure 1: Bitcoin Energy Consumption vs USD Mining Rewards. Source: CBECI33footnotemark: 3and Coinmetrics. [^4] https://ccaf.io/cbnsi/cbeci
44footnotetext: https://coinmetrics.io
Targeted Nakamoto (hereafter ”Targeted”) modifies Nakamoto to create an incentive for miners to push hashrate toward an interval whenever it strays outside the interval. The miner’s minimum block reward is increased when hashrate is below the interval, which induces miners to apply more hashrate. The miner’s maximum block reward is decreased when hashrate is above the interval, which induces miners to apply less hashrate. Monetary neutrality is maintained by offsetting the increase and decrease in the miner’s block reward by opposite adjustments to the spending potential of addresses with UTXO value, in proportion to the value held by each address.
1.1 Two existential risks to Bitcoin
Targeted mitigates two latent, potentially existential threats to Bitcoin. One threat arises from the halving of the minting of UTXO’s (the ”coinbase”), which ensures that transaction fees will eventually become the primary source of block rewards. [^5] Historically, the trend USD/BTC exchange rate has risen sufficiently to offset the halving of the coinbase. However, this cannot continue indefinitely for two reasons. First, coinbase will eventually be zero. Second, maintaining a fixed USD value of the coinbase paid to miners as halving continues would require the inflation adjusted USD to BTC exchange rate to double every four years. This implies a growth rate that exceeds US (and world) GDP by orders of magnitude, which is dynamically unsustainable (Tirole [21]). A transactor is motivated to add a fee to incentivize a miner to include the transaction in its block. [^6] Huberman et.al. [12] A transactor has, at best, a muted incentive to include a fee for network security, which is a non-rival public good that benefits all participants. Historically, transaction fees have comprised less than 10% of total block rewards on average and there is no assurance that transaction fees will increase in the future. [^7] Data on USD block rewards and transaction fees for Bitcoin are compiled by Coinmetrics https://coinmetrics.io The scheduled decline in the coinbase, if not matched by an increase in transaction fees, will cause miners to reduce hashrate, which will lower the cost of an attack on the network. The other threat comes from the reality and the adverse political reaction to the fact that the electricity consumed by Bitcoin miners adds to carbon emissions. Targeted mitigates both of these risks by incentivizing miners to home in on a hashrate interval. This ensures that the cost of attacking the network will not become dangerously cheap while limiting the network’s contribution to carbon emissions and electricity cost. Finally, Targeted preserves aggregate monetary neutrality by proportionally adjusting the spending potential of Bitcoin addresses to offset the minting or burning of UTXO value that that arises from the provision of incentives to miners.
1.2 Links to code and web API
Accompanying this paper is a link to a code-base for controlling Hashrate and a Web API with an empirically estimated hashrate function and a module that enables the user to choose a target hashrate interval and control parameters and to run counterfactual experiments with alternative time-paths of model variables. Both instantiations were developed by Kristian Praizner of MIT.
The code-base is here https://github.com/Krisp140/TargetedNakomoto
The Web API is here https://targetednakamoto.com/About.
Both are anonymized for this submission.
1.3 Limitation
The decrease in network security risk and the increase in electricity cost and carbon emissions as hashrate increases ensures, with a mild restriction on curvature, the existence of a point or interval at which the sum of the direct and externalities costs are minimized. [^8] The restriction is that the function defined as the sum of the costs is is concave. However, locating the minimum requires a numerical assignment of the costs of network security risk and carbon emissions, which lies outside the scope of this paper. A related issue is the tracking of changes in the relationship between hashrate and electricity cost. The efficiency of Bitcoin ASICS has improved over time, reducing the Kw per hash. Coinmetrics [5] estimates that the average electricity per hash among active Bitcoin ASIC’s decreased by half from 2018 to 2022. If improvements continue, a fixed target hashrate interval will be associated with lower electricity consumption and a lower cost of attack. The difficulty here is not the tracking of changes in per kw electricity cost and ASIC efficiency, but rather use of an oracle that it requires, which violates a value held by many in the Bitcoin community. Finally, the relation between electricity cost and security may be non-stationary. A change in the USD value of Bitcoin, or in the size distribution or other characteristics of transactions, may alter the security achieved at any level of hashrate. Against these limitations stand the existential threats mentioned above, which may convince the reader of the need to form consensus on a target interval notwithstanding the gaps in our knowledge.
1.4 Related Literature
This paper builds on the contributions of Nakamoto [17] who laid the foundations for the PoW protocol used in Bitcoin.
There are two previous works with which aspects of this study closely align. Pagnotta [19] identified and modeled the complimentarity between mining hashrate and Bitcoin USD price. The interdependence runs through network security. Hashrate determines security, security affects traders valuation of Bitcoin, which affects it price, price affects the USD value of the mining reward, which determines equilibrium hashrate. Targeted injects into this causal chain a regulator that adjusts the miner’s block reward to incentivize miners to keep hashrate (and thereby security) in the neighborhood of a selected target. Mirkin et.al. [15] show that electricity cost can be reduced, for any level of network security, by adding a ”proof-of-delay” (e.g. using a verifiable delay function) threshold for a miner’s block to be eligible to append to the chain. The mandatory delay interval creates an incentive for miners to shift expenditure from hashrate to investment in mining equipment. The re-allocation of expenditure from variable to fixed cost reduces the carbon footprint. Notably, proof-of-delay does not alter equilibrium expenditure on mining, which is determined by the miner’s block reward. This means it does not affect the cost to attack the network and thereby network security. While proof-of-delay reduces energy consumption for a given level of security, Targeted, which is compatible with proof-of-delay, enables network security to be selected and maintained.
One of the externalities costs addressed in this paper are the carbon emissions generated by the electricity consumed in mining Bitcoin. I rely on two data providers. One is the Bitcoin Electricity Consumption Index (BECI) of the Cambridge U.K. Centre for Alternative Finance [8], which is regarded as a credible - though not infallible - source for information on Bitcoin energy consumption. The other is Coinmeterics  [^9] https://coinmetrics.io(which is a source for BECI) for other Bitcoin performance indicators. Coinmetrics is considered to be a careful aggregator of data on cryptocurrencies. Sedlmeir et.al. [20] and Gallersdorfer et.al. [10] review comparative magnitudes of energy consumption for different blockchain protocols. Gallersdorfer [10] compares the estimation of Bitcoin electricity consumption across several sources. For my purposes, the salient finding is that these studies document a time-path of energy consumption that is tightly correlated with the time-path of hashrate in Bitcoin.  [^10] See Figure 1 in Gellersdorfer [10].
As with several previous studies of blockchain protocol, this paper uses mechanism design approach to the study of the consensus protocol. Chen et.al. [14] surveys various aspects of existing PoW blockchain protocols, including the incentive to increase energy consumption in mining, from a mechanism design perspective. Cong et.al [6] analyses the effect of PoW mining competition on energy consumption. Leshno and Strack [13] identifies features of the miner payoff structure that are required to maintain blockchain stability. They prove that certain of the core properties of a decentralized PoW blockchain - namely free entry of anonymous profit driven miners - are only incentive compatible with a reward mechanism that is similar to Nakamoto. Targeted preserves those features.
In their study of the Bitcoin payment system, Huberman et.al. [12] characterize an equilibrium where transactions are appended to blocks in descending order of the size of fee offered by transaction senders, which reflects transactor’s relative willingness to pay for a position in the queue for placement on the blockchain. The adjustment to the miner’s block reward in Targeted is guided by this insight. When the block reward is adjusted, the ordering of the rewards received by the miner from transactions always matches the order of fees offered by transaction senders. This, in turn, leaves unaltered he order in which transactions are appended to blocks.
1.5 The research question and design constraints
Targeted augments Nakamoto by creating an incentive for miners to home in on a hashrate interval. The augmentation comes close to adhering to formal constraints which reflect core values held by the Bitcoin community. [^11] What constitutes a core feature is subjective. I developed the list herein after consultation with a number of participants in the Bitcoin network, including developers, miners and exchange operators. Ultimately, the reader will judge for herself whether my list is complete. The constraints are partitioned into the consensus protocol layer (C constraints) and the transaction protocol layer (T- constraints).
[…]
It will be shown that Targeted satisfies each of these constraints with one exception. Targeted can instruct miners to burn UTXO’s from transaction outputs, which may require a hard-fork to implement.
1.6 Roadmap
The rest of the paper is organized as follows. Section 2 provides an overview of the two types of cost, the sum of which Targeted is designed to minimize; costs that increase with hashrate and those that decrease with hashrate. Section 3 casts Nakamoto and Targeted in a mechanism design framework in order to articulate the key features of each protocol and their interaction and to show how Targeted augments Nakamoto. Section 4 shows how the ratio of mining puzzle difficulty to blockchain growth rate is a sufficient statistic for hashrate and then derives the expression for equilibrium hashrate as a function of, inter alia, the value of the block reward. Section 5 uses the relationship between the puzzle difficulty/growth rate ratio and hashrate, along with hashrate equilibrium to construct the policy for adjusting the miner’s block reward to incentivize miners to push hashrate toward the target hashrate interval. Section 6 demonstrates that the hashrate control algorithm dynamically pushes hashrate toward the target interval. Section 7 states a protocol that renders the aggregate money supply and the distribution of spending potential unaffected by the Hashrate Control Algorithm. The paper concludes with a summary of the key properties of Targeted and call attention (again) to the urgency for the Bitcoin community to address the issues discussed in this paper.
2 Externalities Costs
This section explains the costs associated with honest miner hashrate. Section 2.1 reviews evidence of a link between Bitcoin hashrate and carbon emissions. Section 2.2 discusses three vectors of attack on the network that become more feasible as hashrate decreases; double-spend, eclipse and shorting attacks. In Section 2.3 derives a target hashrate interval as a function of externalities costs.
2.1 Externalities costs that increase with hashrate
Hashing, or puzzle guessing, involves computation that consumes electricity. This generates two external effects. One is to increase the price of electricity and the other is to increase the social cost of carbon emissions from the from fossil fuel burning electricity generators. [^12] Chen et.al. [14] model the flow of carbon emissions from multiple energy sources. These costs are related in the sense that an increase in hashrate will, by increasing electricity demand, put upward pressure on electricity price and, to the extent that the increase in demand is met by an increase in electricity generation (i.e. that electricity supply is not inelastic), cause a rise in carbon emissions. [^13] Fowlie et.al. [9] estimate a price elasticity of demand for electricity in the U.S. of -2.06 for with a range of alternative elasticities around -1 (see Table 2). These results show that an increase in electricity demand will be split between increases in both price and quantity.
2.2 Externalities costs that decrease with hashrate
We review three vectors of attack on a PoW blockchain that decrease in profitability as hashrate increases.
Double-spend attack A double-spend attack requires the attacker to build an alternative chain that overtakes the incumbent canonical chain at in terms of cumulative mining puzzle difficulty. An attacker with over 50% of total hashrate will overtake the incumbent canonical chain at some block with probability measure 1. At less than 50% share, the attacker’s probability of success is less than half. [^14] See Gervais et.al. [11] for simulations of attacks using different ratios of attacker hashrate to miner hashrate. The frequency of success is increasing in attacker hashrate and the average number of blocks required to overtake the incumbent canonical chain is decreasing in attacker hashrate. The amount of attacker hashrate required to successfully fork the blockchain, and its cost, is increasing in the level of honest miner hashrate. Budish (2018) [3] and Aronoff et.al. [1] point out that the cost of a successful attack is at least partially offset by the block rewards that accrue to the attacker. The possibility that a victim of a double-spend attack retaliates by mining on the incumbent canonical chain leads to a war of attrition between the attacker and the victim (Moroz et.al. [16]). Aronoff et.al. [1]) point out that the outcome of a war of attrition depends on equilibrium refinements and exigent circumstances, even when the costs between players differ. Nevertheless, an increase in the cost of launching the attack will act as a deterrent by increasing the sunk cost that the attacker loses if it is unsuccessful. Equation 1 is the attacker’s un-discounted ex-ante expected profit when block rewards equal hashrate cost. The key observation is that, for any probability of success, the attacker’s expected profit is declining in honest miner hashrate.
Ex-ante Expected $\text{Ex-ante Expected Profit}\;=\;Prob(success)V-\\ \{1-Prob(success)\}\text{E[hashrate cost per block]E[number of blocks]}$ (1)
Where $V$ is the double-spend value.
Eclipse attack An eclipse attack involves isolating a node from the network and constructing a chain, forked from the incumbent canonical chain, that is broadcast only to the victim (the ”false chain”). On the false chain the attacker sends the victim UTXO value, which the victim believes to be on the canonical chain. After the attacker receives the exchange item from the victim, it discontinues applying hashrate to the false chain. The attacker incurs two elements of cost. One element is the hashrate the attacker applies to the false chain. The other element is the cost of isolating the victim from the network. Neither of these costs necessarily increase with honest miner hashrate. However, the attacker may need to apply hashrate that is proximate to honest miner hashrate to convince the victim into believing that the fake chain is canonical. Therefore it is possible that the cost of an eclipse attack is increasing in honest miner hashrate, in which event the profit is declining in honest miner hashrate.
Shorting attack A shorting attack has two parts; (i) placing a bet that a cryptocurrency will decline in value relative to some other currency or asset and (ii) gaining control of the cryptocurrency network in order to cause a disruption that will induce a loss in confidence and fall in exchange value. One type of bet is a forward sale of, say, Bitcoin in dollars. The attacker contracts to sell BTC to a counterparty at a future date for a dollar strike price per unit of BTC. The attacker anticipates that its disruption to Bitcoin will enable it to purchase BTC at a dollar price that is below the strike price, enabling it to resell to the counterparty at a profit. The size of the bet is independent of hashrate. In a liquid futures market an attacker could contract to purchase \$x dollars worth of BTC at some future price regardless of the level of hashrate. On the other hand, the cost of gaining control of the network scales with hashrate. Thus, the profitability of the attack is declining in hashrate. [^15] The value of the block rewards earned by the attacker decline with the Bitcoin exchange value.
2.3 The target hashrate interval
Figure 2 is a conceptual depiction of the curves representing network security and energy externalities cost. The horizontal axis is the hashrate of miners in the Bitcoin network. The vertical axis represents the cost of security and energy externalities. Total cost is the summation of these curves at each level of hashrate. The goal of Targeted is minimize the total cost of security and energy externalities. The negative slope of the network security curve reflects that the network becomes less vulnerable as hashrate increases. Network security vanishes as hashrate approaches zero, since the cost of acquiring 51% hashrate is negligible (equivalent to unbounded security cost). The cost of attacking the network becomes extremely expensive at very high levels (equivalent to negligible security cost). A mild smoothness assumption implies that the network security curve is convex in hashrate. I assume the unit hashrate cost is either constant or increases with electricity consumption (and hashrate). The shape of the social cost of carbon emissions is convex in electricity consumption (and hashrate).161616 ”The best evidence is that impacts of climate change will be non-linear and cumulative.” (William Nordhaus Nobel Prize lecture in economics [18]). Deitz et.al. [7] identify the ”three essential elements of the climate problem – the endogeneity of growth, the convexity of damage and climate risk”. The energy externality cost is the summation of these two costs, which is convex. The sum of the two curves is convex, which assures the existence of an interval in which total cost is minimized. A reduction in per hash enery cost induced by improvements in ASIC efficiency and/or electricity cost will shift the network security curve to the right. A reduction in per hash carbon emissions from an increase in the use of clean energy will shift the hashrate/energy externalities cost curve to the right.
Figure 2: The Target Hashrate Interval
3 Targeted Nakamoto - A Mechanism Design Perspective
This section describes the elements of Targeted and compares the mechanism designs of Targeted and Nakamoto.
3.1 Key Building Blocks of Targeted Nakamoto
Targeted controls hashrate by leveraging four features that are intrinsic to the blockchain. One feature is informational. The hashrate, which is not recorded on the blockchain, can be estimated from mining puzzle difficulty and block timestamps, which are recorded on the blockchain. The augmented protocol targets bounds on hashrate, for which the ratio of puzzle difficulty to blockchain growth rate is a sufficient statistic.
The second feature is an object that can be modified. The Nakamoto block reward is comprised of two parts; an algorithmic minting of new tokens and a discretionary fee added to a transaction. Both parts of the block reward are recorded in the blockchain and can be altered by the code. Targeted places a floor or ceiling on the the miner’s block reward when hashrate is above or below the target interval.
The third feature is a market incentive. The hashrate applied by miners is affected by the block reward; it increases as the value of the block reward increases and decreases as the value of the block reward decreases. This implies that hashrate can be controlled by manipulation of the block reward.
The fourth feature enforces monetary neutrality by adjusting the aggregate UTXO spending potential to exactly offset the amount added to or subtracted from the total block reward to obtain the miner’s block reward.
3.2 An overview of the mechanism design of the protocols
Targeted can be decomposed into two protocols that interact while operating separately. The Nakamoto part is the mechanism associated with the current protocol. The Targeted part adds a mechanism to achieve a target hahsrate interval. Figure 3 displays the information recorded on the blockchain that is used by each protocol [^17] The network also records the addresses of UTXO’a at each block. Neither Nakamoto or Targeted use that information..
Figure 3: Blockchain Records
Table 1 shows the basic elements of Nakamoto and Targeted. For each protocol, the information contained in the signal, the instrument and the feedback are recorded on the blockchain. There is a salient difference in the target. The target for Nakamoto, blockchain growth rate, is recorded on the blockchain. The target for Targeted, hashrate, is not recorded on the blockchain.
Protocol Mechanism Design
Element Nakamoto Targeted
Target Blockchain Growth Rate Hashrate Target Interval
Signal Block Timestamps Puzzle Difficulty/Growth Rate
Feedback Block Timestamps Puzzle Difficulty/Growth Rate
Instrument Puzzle Difficulty Block Reward
Policy Adjust to Growth Rate Ceiling or Floor on Block Reward
Table 1: A Comparison of Mechanisms
4 The Model
This section derives and characterizes the equilibrium level of hashrate. Section 4.1 derives the function of puzzle difficulty that is a sufficient statistic for hashrate. Section 4.2 characterizes the equilibrium hashrate as a function of, inter alia, the miner’s block reward.
4.1 The puzzle difficulty signal of hashrate
Nakamoto sets a target time for intervals between the addition of new blocks to the blockchain. The timestamps are read by the code at intervals of $m$ blocks, which is an epoch. An epoch in Bitcoin $m=1,260$ blocks. The instrument used to achieve the target is an adjustment to the difficulty of the mining puzzle, denoted as $D$. The mining puzzle is the solved when the miner’s input into a hash function outputs a value that falls within the target range set by the code. Puzzle difficulty is the ratio of the range set by the code to the range of all possible output values. The probability of guessing the solution is $p=1/D$ and each guess is independent. [^18] By construction $p$ is a monotone decreasing function $P(D)$ of $D$. The derivative $\frac{dp}{dD}<0$ in both the simplified expression in the text as well as e.g. the algorithm used in Bitcoin Core. The derivative is the only property material to our analysis. $D$ is adjusted after the $m$th (final) block of a epoch by dividing the number of blocks into the elapsed time from the previous timestamp. When the blockchain growth rate exceeds the target, puzzle difficulty is increased. When the blockchain growth rate falls below target, puzzle difficulty is decreased.
A single mining computer can guess a puzzle solution sequentially in time. [^19] A computer with $n$ parallel processors is treated here as $n$ separate computers. I model mining computers as making successive guesses synchronously at times $t\in\{t_{1},t_{2},...\}$, each separated by an interval of temporal length $t_{i+1}-t_{i}$ which I normalize to 1 unit of time [^20] The synchronicity enables the expression of the time-path of puzzle solutions compactly as a Bernoulli Process, which simplifies the analysis WLOG.. The Nakamoto blockchain growth target is one block every $T^{*}$ units of time. During this time interval each computer has guessed (or hashed) $T^{*}$ proposed solutions. I define a unit of hashrate as a single mining computer making $T^{*}$ puzzle guesses. If there are $N$ mining computers making guesses, the number of total guesses made in the time interval of the blockchain growth target is $NT^{*}$ and the hashrate is $N$212121 This accords with the conventional definition of hashrate used e.g. by Coinmetrics [4], as puzzle guesses per second. The framework for guessing puzzle solutions over time is a Bernoulli Process. Let $T$ denote the time the first puzzle solution is broadcast to the network. For a hashrate of $N$ the expected time to reach the solution is expressed by Equation 2.
[…]
Equation 2 shows that, in order to achieve the growth rate target of $T^{*}$ in expectation, puzzle difficulty $D$ must be adjusted to offset hashrate $N$. Since $T$ for the prior epoch is recorded on the blockchain and $D$ is set by the code, the formula can be inverted to estimate hashrate as a function of the growth rate and puzzle difficulty, $\hat{N}=D/T$232323 $\hat{N}=N+\epsilon$, where $\epsilon$ is a mean zero iid random variable. $\hat{N}$ is unbiased since the observed $T$ is an unbiased estimate of the mean time interval implied by the Bernoulli Process. It is also a sufficient statistic for hashrate since $D$ and $T$ contain all of the information on $N$ that is encoded on the blockchain.
4.2 Mining equilibrium
The miner who is first to broadcast a puzzle solution receives a block reward of cryptocurrency. Miners incur cost in terms of the local official currency, which we denote USD. [^24] Profit can be expressed in terms of cryptocurrency. In that case the unit hashrate cost cost $c$ is multiplied by $1/e$, which is the USD to BTC exchange rate. The equilibrium hashrate is the same whatever currency is chosen to be the unit of account. Mining equilibrium is a state where miners choose hashrate to maximize expected profit. When miners adjust hashrate at short block intervals profit maximization implies that the profit from a marginal unit of hashrate equals the competitive rate of profit. The independence of puzzle guesses means each unit of hashrate has a $1/N$ probability of receiving the block reward. Equation 3 is the equilibrium profit from mining one unit of hashrate (i.e. operating one mining computer for $T^{*}$ units of time). [^25] Equation 3 is a dynamic equilibrium in the sense that it includes a growth rate that deviates from target. A rigorous expression would show $\Delta_{T}$ as a function of puzzle difficulty and hashrate. We do not derive the equilibrium from fundamentals and therefore we do not prove the stability of the solution.
$\pi=\frac{1}{N}eP_{B_{n}}^{M}(1+\Delta_{T})-c=\phi$ (3)
$\Delta_{T}=T^{*}-T$ is the difference between the target growth rate, $T^{*}$, and the expected actual growth rate $T$ [^26] Blockchain growth rate is expressed as an expectation since it is stochastic. $(1+\Delta_{T})$ is the number of blocks per $T^{*}$ units of time; $c$ is the marginal cost of a unit of hashrate (i.e. one computer for $T^{*}$ units of time); [^27] Marginal cost is the cost of electricity to run the ASIC’s. Fixed costs, such as the rent on the building where the computers are stored, or the sunk cost of purchasing the computers, are not included. $e$ is the USD/BTC exchange rate and $p_{B_{n}}^{M}$ is the expected miner’s block reward on the $n^{th}$ block. ($p_{B_{n}}$ is the total block reward). I make no assumption about the structure of the mining market. $\phi\geq 0$ a competition parameter. Under free entry in the mining market $\phi=0$. Under less competitive conditions $\phi>0$. Equation 4 is the mining market equilibrium.
$N=eP_{B_{n}}^{M}(1+\Delta_{T})/(c+\phi)$ (4)
Equation 4 shows that hashrate $N$ is a function of the market conditions faced by miners. Puzzle difficulty does not appear in Equation 4. Equation 5 substitutes $D/T$ for $N$ in Equation 4.
$D/T=eP_{B_{n}}^{M}(1+\Delta_{T})/(c+\phi)-\epsilon$ (5)
Equation 5 displays the monotone relationship between the miner’s block reward, $p^{M}_{B}$, and the sufficient statistic for hashrate, $D/T$. The dynamic adjustment of hashrate to changes in the block reward is evaluated in Section 6.
5 The Targeted Block Reward Policy
This Section explains the implementation of the policy of adjusting the miner’s hashrate. Section 5.1 shows how the allocation of an addition or subtraction to the miner’s block reward leaves unchanged the miner’s incentive to append transactions onto blocks according to the ranking of the size of fee offered by the transactor. Section 5.2 builds on the analysis of Section 4 to state the policy of adjusting the miner’s block reward in response to the ratio of puzzle difficulty to blockchain growth rate. Section 5.3 states the algorithm for controlling the miner’s block reward.
5.1 The block reward adjustment allocations
One of the T constraints is that $Targeted$ should not change the order in which transactions are appended to blocks. Huberman et.al. [12] demonstrate that Nakamoto induces an equilibrium ordering of transaction fees and transactions on the blockchain that reflects transactors’ relative time preferences. Transactions are appended to blocks in descending order of the size fee offered and the ranking of fees match the ranking of transactors’ time-preference for executing their transaction. [^29] See Theorem 1 and Proposition 2. Note that the result relies on parametric assumptions. The reader can consult Huberman et.al. [12] for details. It is a constrained efficient outcome in the sense that if agents A and B send their respective transactions to the mempool at the same time and agent A incurs a higher cost for delay in completing its transaction than agent B, agent A’s transaction will be appended to the blockchain ahead of agent B’s transaction. Targeted preserves the same ordering of transactions by adjusting the coinbase and each transaction fee by the same percentage. $Fee_{f,B_{n}}$ denotes the fee paid by the $f$th transaction appended to block $B_{n}$. Denoting $\xi=(p^{M}_{B_{n}}-p_{B_{n}})/p_{B_{n}}$, the ordering $Fee_{i,B_{n}}>Fee_{j,B_{n}}$ implies the ordering $\xi Fee_{i,B_{n}}>\xi Fee_{j,B_{n}}$, i.e. the ordering of transaction fees sent to the miner is unchanged. $R_{n}$ denotes the coinbase portion of the block reward. The miner’s block reward is $p^{M}_{B_{n}}=(1+\xi)\big{[}\sum_{f}Fee_{f,B_{n}}+R_{n}\big{]}$. When $\xi<1$ the quantity $\xi p_{B_{n}}$ UTXO’s is burned. When $\xi>1$ the quantity $\xi p_{B_{n}}$ UTXO’s are minted (above the scheduled coinbase).
Crucially, this method of proportional adjustment of the fee paid to the miner does not alter the Nash equilibrium of the pattern of fees offered by transactors. This is readily apparent by considering the incentive for $A$ or $B$ to alter their fee after a proportional adjustment $\xi$ of the miner’s fee. Miners will still have an incentive to append transactions to blocks in ascending order of fees. If $B$ does not change its fee, $A$’s incentives are unaffected. It still must offer a higher fee than $B$ to secure a position in the queue of transactions ahead of $B$. Similarly, if $B$ was unwilling to offer a higher fee than $A$ before the adjustment, it not have an incentive to do so after the adjustment.
5.1.1 Addition and Subtraction of UTXO Value
When $p^{M}_{B_{n}}-p_{B_{n}}>0$, the difference is reflected in an additional payment to the miner above the total block reward. When $p^{M}_{B_{n}}-p_{B_{n}}<0$, the difference is reflected in a subtraction in payment to the miner from the total block reward. The source of increased UTXO value in the former case (minting or sending from another address) and the allocation of the unused UTXO value in the latter case, are dealt with by the Targeted monetary policy in Section 7.
5.2 The puzzle difficulty signal and policy switch points
Equation 4 shows that hashrate $N$ is over-determined. In addition to the miner block reward $P_{B}^{M}$, it is affected by mining cost $c$ and the exchange rate $e$. These variables move independently of each other. This means the block reward instrument will not be able to perfectly achieve the target. To compensate, $Target$ dynamically adjusts the block reward to move hashrate toward the target interval over time.
The mechanism works as follows. The endpoints of the hashrate target interval, $\{N_{UB},N_{LB}\}$ map into puzzle difficulty signals $\{\frac{D}{T}(N_{UB}),\frac{D}{T}(N_{LB})\}$, which bound the puzzle difficulty target interval. An adjustment to the miner’s block reward can be made after the end of each epoch when puzzle difficulty is adjusted. Epochs are numbered sequentially in temporal order, starting from 1. Epoch 1 includes blocks numbered 1 (Genesis) to $m$. Epoch 2 includes blocks numbered $m+1$ to $2m$, and so forth. The end-block of epoch $q$ is denoted $E_{q}$. The ratio $D/T$ is measured for epoch $q$ at block $E_{q}$ (the ”adjustment block”). $D$ is the puzzle difficulty in effect for epoch $q$ and $N$ is the average hashrate in epoch $q$. It does not include the adjustment to $D$ made after $E_{q}$ is appended to the blockchain. The monotone relationship between the sufficient statistic and hashrate implies that hashrate is above target when $D/T>\frac{D}{T}(N_{UB})$ and below target when $D/T<\frac{D}{T}(N_{LB})$.
Targeted imposes a ceiling on the miner’s block reward in epoch $q+1$ when $D/T$, measured for epoch $q$, is above the target interval. After a ceiling has been put in place, it is lowered at each subsequent adjustment block until $D/T$ drops below $\frac{D}{T}(N_{UB})$. Similarly, Targeted imposes a floor on the miner’s block reward in epoch $q+1$ when $D/T$, measured for epoch $q$ when $D/T$ is below $\frac{D}{T}(N_{LB})$. When a floor has been put in place, it is raised at each subsequent adjustment block $E_{q}$ until the sufficient statistic rises above $\frac{D}{T}(N_{LB})$. When $D/T$ is inside the target interval, if a ceiling is in place from the prior adjustment block, the ceiling is raised, and so forth until it is removed altogether. If a floor is in place from the previous adjustment block, the floor is lowered at the current adjustment block, and so forth until it is removed altogether. The switching is displayed in Figure 4. At $E_{1}$ the sufficient statistic crosses the $\frac{D}{T}(N^{UB})$ switch-point from below. This turns ”on” the instrument that imposes a ceiling on the miner block reward, which is lowered at epoch $E_{2}$ and raised at $E_{3}$. $D/T$ is below $\frac{D}{T}(N_{LB})$ at $E_{4}$, which causes the ceiling to be removed and a floor to be imposed. The floor is raised at $E_{5}$.
Figure 4: Mining Puzzle Difficulty Signal Switch Points
The dynamic feature of the policy can be seen when another RHS variable of Equation 5 is increasing. For example, if the exchange rate is rising when $D/T$ is above $\frac{D}{T}(N_{UB})$, the block reward ceiling will be pushed down at each adjustment point. It may transpire that hashrate continues to rise for some period of time, but so long as the exchange rate is bounded, the absolute value of the block reward ceiling adjustments will eventually rise above the exchange rate increase, which will push down hashrate.
5.3 Policy to control hashrate
The Hashrate control algorithm controls the imposition of ceilings and floors on the miner’s block reward. [^30] Appendix A contains Python code for the Hashrate Control Algorithm.
The Hashrate Control Algorithm A ceiling or floor on the miner’s block reward is adjusted after the end-block $E_{q}$ of each epoch $q$ as a function of $D/T$. There are four cases to consider.
Set parameters $\tau,\gamma\in[0,1]$.
Case 1: Hashrate is above the upper bound of the hashrate interval, $\frac{D}{T}(N_{UB})$.
A ceiling on the miner’s block reward is imposed. For each block in epoch $q+1$, payment to miner is capped at $\text{ceil}_{q}=\tau\overline{P_{B}}(q),\;\tau<1$, where $\overline{P_{B}}(q)$ is the median block reward in epoch $q$.
For each subsequent epoch until $\frac{D}{T}(N_{UB})$ is crossed from above, the ceiling is lowered by the formula $\text{ceil}_{q+1}\leftarrow\tau\text{ceil}_{q}$.
Case 2: The puzzle adjustment at block $E_{q}$ implies that $D/T$ is below the lower bound of the target hashrate interval, $\frac{D}{T}(N_{LB})$.
A floor on the miner’s block reward is imposed. For each block in epoch $q+1$, payment to miner is capped at $\text{floor}_{q}=(1+(1-\tau))\overline{P_{B}}(q)$, where $\overline{P_{B}}(q)$ is the median block reward in Epoch $q$.
For each subsequent epoch until $\frac{D}{T}(N_{LB})$ is crossed from below, the floor is raised by the formula $\text{floor}_{q+1}\leftarrow(1+(1-\tau))\text{ceil}_{q}$.
Case 3: The puzzle adjustment at block $E_{q}$ implies that $D/T$ is inside the target hashrate interval.
For each subsequent epoch until $\text{ceil}_{q}\leq\overline{P_{B}}(q)$, the ceiling is increased by the formula $\text{ceil}_{q+1}\leftarrow(1+(1-\gamma)\text{ceil}_{q}$ Thereafter, the ceiling is removed.
For each subsequent epoch until $\text{floor}_{q}\geq\overline{P_{B}}(q)$, the floor is reduced by the formula $\text{floor}_{q+1}\leftarrow\gamma\text{floor}_{q}$ Thereafter, the floor is removed.
 [^31] The formulas $\tau$ and $(1+(1-\tau))$ ensure symmetry in the adjustment of the ceiling and the floor, i.e. that the ceiling and floor adjust by the same percentage each epoch.
Finally, note that the Hashrate Control Algorithm is a hard-fork, since the imposition of the floor to the miner’s block reward can require the minting of coinbase that exceeds the amount scheduled in Nakamoto.
6 Dynamic Adjustment of Hashrate
This section evaluates the dynamic response of hashrate to changes in the value of the miner’s block reward which occur when hashrate is outside the target interval. The dynamic adjustment of hashrate to a change in the miner’s block reward can be evaluated when hashrate adjusts to maintain a profit $\phi$ per target time $T^{*}$. [^32] The assumption of a constant profit rules out the possibility of a dynamic strategy to influence hashrate by a miner coalition. Arguably, $\phi$ internalizes the exercise of market power.
Consider an epoch $q$ where the total block reward is constant at $p^{M}_{B}$ and the puzzle difficulty adjustment at the end of epoch $q-1$ achieves the target growth rate when hashrate is $N$, and $N$ is the equilibrium hashrate when the miner’s block reward is $p^{M}_{B}$. Suppose a ceiling or floor is imposed at the end of epoch $q-1$ resulting in a marginal change in the miner’s block reward equal to $\Delta p$ (i.e. holding fixed the exchange rate, $e$, and hashrate cost $c$). Let $\Delta_{T}$ denote the distance between the actual blockchain growth rate and the target blockchain growth rate induced by a change in hashrate in epoch $q$, denoted $\Delta N$.
$\Delta_{T}=T^{*}-T=\frac{D}{N}-\frac{D}{N+\Delta N}$
Equation 6 is the equilibrium of an interval of $T^{*}$ units of time in epoch $q$.
[…]
The change in hashrate can be decomposed into a lottery effect and a growth rate effect. The lottery effect reflects that a change in hashrate changes the odds of any unit of hashrate mining the next block. The growth rate effect reflects that a change in hashrate alters the blockchain growth rate, which changes the number of puzzle lotteries in which a unit of hashrate can compete in $T^{*}$ units of time. Noting that $\frac{D}{N}|modulo\;T^{*}=1$, the growth rate effect is $2-\frac{D}{N+\Delta N}$. Equation 7 is the marginal effect of hashrate on profit.
$\frac{\Delta\pi}{\Delta N}=\big{(}e(P_{B}^{M}+\Delta p)\big{)}\Big{[}\frac{\Delta\text{lottery effect}}{\Delta N}+\frac{\Delta\text{growth rate effect}}{\Delta N}\Big{]}\\ =-2e(P_{B}^{M}+\Delta p)\frac{1}{(N+\Delta N)^{2}}<0$ (7)
Equation 8 is the marginal effect of the miner’s block reward on profit.
$\frac{\Delta\pi}{\Delta p}=\text{lottery effect}\times\text{growth rate effect}\times p^{M}_{B}>0$ (8)
[…]
Proposition 1.
The marginal effect of the Hashrate Control Algorithm is to dynamically push hashrate toward the target interval.
7 Monetary Neutrality
This section describes a protocol that ensures the implementation of the Hashrate Control Algorithm does not cause any substantive change to the Nakamoto monetary policy in terms of the spending potential of addresses. Targeted induces offsetting additions or subtractions from the UTXO value of an address that occurs when the address sends its UTXO value in a transaction. The concept of money supply is adapted to include the UTXO adjustments that are committed for future transactions.
Definition 1.
Spending Potential
(i) Address Spending Potential The UTXO value in address $x$ plus the UTXO value that is scheduled to be added to or subtracted from $x$ when it sends a transaction.
(ii) Aggregate Spending Potential The sum of the spending potential of addresses plus the increase or decrease in spending potential that is not currently assigned to addresses, but which under the protocol can be assigned in the future.
The concept of monetary neutrality has two dimensions.
Definition 2.
Monetary Neutrality
(i) Aggregate Monetary Neutrality Aggregate UTXO spending potential under Targeted is the same as Nakamoto at every block.
(ii) Distributional Monetary Neutrality At any block, an addition or subtraction of UTXO from the total block reward is allocated to an address in proportion to its percentage of aggregate spending potential.
7.1 The Targeted Monetary Policy
This section presents a high-level description of the monetary policy as it affects a single address $x$. Python code describing the logic of the policy is in Appendix B.
Subtractions from the total block reward: When a UTXO value of $q$ is subtracted from the total block reward at block $B_{n}$ the miner is instructed to send $q$ to an address denoted the ”UTXO pool”. The UTXO pool is a single address from which UTXO value is sent to other addresses and has an initial value $P$ to which $q$ is added. Every address $x$ holding UTXO value is assigned spending potential in the UTXO pool (the ”assigned value”) which is denoted $e_{x}$. When $q$ is sent to the UTXO pool, $x$’s assigned value is increased by $x$’s fraction of aggregate UTXO spending potential implied by the state after the block $B_{n}$ transactions, denoted $S_{x|B_{n-1}}$ ($x$’s ”percentage”) times $q$, rounded down to the nearest Satoshi.
Additions to the total block reward: When $q$ is added to the total block reward at block $B_{n}$ there are two possible adjustments to accounts. If $P-q>0$, $x$’s assigned value in the UTXO pool is reduced by the product of its percentage times $q$, rounded down to the nearest Satoshi. When $P-q\leq 0$ two accounting changes occur; (i) P and $x$’s assigned value $e_{x}$ are each reduced to zero and (ii) the value $q-P$ is added to an account that tracks value that is subtracted from the spending potential of addresses (the ”Remainder Accounts”) which has initial value $R$. Each address $x$ is assigned a value in the remainder account, denoted $r_{x}$, which is increased by its percentage times $P-q$, rounded down to the nearest Satoshi. [^37] The UTXO value subtracted from P can either be burned or sent to the miner. The latter would reduce the minting of UTXO required to increase the miner’s block reward.
Note that the spending potential of an address $x$ with UTXO value $x$ is $x+e_{x}-r_{x}$.
Rounding Adjustment At each block $B_{n}$ the unassigned UTXO value in the UTXO pool and the Remainder Accounts are assigned to the individual address accounts, $e_{x}$ and $r_{x}$ in accordance to the percentage of address $x$, rounded down to the nearest Satoshi.
Transactions Prior to the adjustments to the UTXO pool and Remainder pool, the transactions in block $B_{n}$ are executed. An adjustment is made when address $x$ sends UTXO value in a transaction that is appended to block $B_{n}$. If $e_{x}-r_{x}>0$, the difference is sent to $x$. If $r_{x}-e_{x}>0$, the difference is burned from $x$’s UTXO value. [^38] The burning requirement invalidates the transaction when the burn value exceeds the unspent UTXO. The initial values of $e_{x}$ and $r_{x}$ are the ending values for block $B_{n-1}$. The output values of $S_{x|B_{n,1}},e_{x}$ and $r_{x}$ are the initial state variables in the UTXO pool and the Remainder Accounts adjustments.
Figure 5 depicts the operation of the Targeted monetary policy at block $B_{n}$. The timing sequence is denoted by the circled numbers. The first step is the execution of the transactions. The adjustment to a transaction sender’s UTXO value is determined by the difference between its account value in the UTXO pool and its account value in the Remainder Accounts. The second step is the payment of the miner block reward. If $P^{M}_{B_{n}}<p^{M}_{B_{n}}$, UTXO value of $q$ is minted to cover the payment. The third step is to send of UTXO value $q$ to the UTXO Pool if $P^{M}_{B_{n}}<p^{M}_{B_{n}}$ or else burn UTXO value up to $q$ in the UTXO pool and increase the value of the Remainder Accounts by the residual. Adjustments to the address account values in the UTXO Pool and the Remainder Accounts are made.393939 Steps 2 and 3 can be simultaneous. When $q<0$ the minting of UTXO value is reduced by the amount of value in the UTXO pool, which is sent to the miner from the UTXO pool. This is implicit in the code in Appendix B.
Figure 5: Targeted Monetary Policy
Proposition 2.
Targeted Achieves Monetary Neutrality
Proof.
The difference between the total block reward and the miner’s block reward, $q$, is offset by changes of opposite sign in the value of the UTXO pool and the Remainder Accounts; $q=\Delta P+\Delta R$. This proves aggregate monetary neutrality. The value change in the UTXO Pool and the Remainder Accounts are allocated to addresses in proportion to their spending potential, rounded down to the nearest Satoshi and the unassigned residual is assigned to addresses in accordance with the same formula at each block. This proves distributional neutrality. ∎
8 Conclusion
Targeted is motivated by two observations about PoW blockchains. One observation is that there are costs which increase with hashrate - electricity and carbon emissions - and costs which decrease with hashrate - network security. This observation suggests there is an interval of hashrate in which total costs are minimized. The other observation is that hashrate is a function of the value of the block reward, which is unbounded. The Nakamoto protocol does not provide an incentive for miners to avoid hashrate dropping to a very low level or reaching a very high level. Targeted incentivizes miners to home in on a target hashrate interval. This is achieved by imposing a floor to the miner’s block reward when hashrate is below target and imposing a ceiling on the miner’s block reward when hashrate is above target. The ceiling is progressively lowered and the floor is progressively raised for so long as hashrate is outside the target interval. Targeted achieves aggregate and distributional monetary neutrality. Aggregate monetary neutrality means that the UTXO spending potential of the network is the same under Targeted and Nakamoto at each block. Distributional monetary neutrality means that the UTXO spending potential of addresses are adjusted to offset increases and decreases in the block reward in proportion to their UTXO value.
There are two preconditions for implementing Targeted. One is that a target electricity consumption interval needs to be identified. The second is that the efficiency of mining computers needs to be estimated and recalibrated over time in order to adjust the target hashrate interval to ensure it maps into the target electricity consumption interval. For example, when ASIC’s become more efficient, the target hashrate interval will need to be adjusted upward, since each hash consumes less electricity.
Finally, I argued that there is an urgency to augment the Nakamoto protocol to control hashrate. The current perceived high level of electricity consumed by Bitcoin miners has contributed global warming and a political backlash against Bitcoin. This is a near term challenge for Bitcoin. In the long run Bitcoin faces an existential threat from the opposite direction. The scheduled halving of coinbase approximately every four years will eventually make transaction fees the primary source of miner revenue. Historically, transaction fees have comprised a small percentage of overall block reward. Therefore, in the absence of an orders of magnitude increase in transaction fees, hashrate and electricity consumption may decline to a low level that threatens network security.
Appendix A Code for The Hashrate Control Algorithm
This appendix states the logic steps of the Hashrate Control Algorithm. It is not the code that would be written to implement the Targeted on servers in the Bitcoin network.
 1: Python code for Hashrate Control Algorithm
class Epoch:
def __init__(self):
self.rewards = [] # To store the rewards of each block
within an epoch
self.hashrates = [] # To store the hashrates of each block
within an epoch
self.ceil = None
self.floor = None
self.difficulty = None
self.elapsed_time = None # T_n
def median_block_reward(self):
sorted_rewards = sorted(self.rewards)
return np.median(sorted_rewards)
def average_hashrate(self):
sorted_hashrates = sorted(self.hashrates)
return np.mean(sorted_hashrates)
def add_reward(self, reward):
self.rewards.append(reward)
def add_hashrate(self, n):
self.hashrates.append(n)
class Blockchain:
def __init__(self, tau, gamma, upper_bound, lower_bound, initial_difficulty):
self.tau = tau
self.gamma = gamma
self.epochs = [Epoch()]
self.DT = None # Puzzle difficulty/Blockchain growth rate;
get this value from blockchain source code
self.DT_N_UB = upper_bound # Upper boundary
self.DT_N_LB = lower_bound # Lower boundary
def adjust_reward(self, reward, epoch_idx):
epoch = self.epochs[epoch_idx]
if epoch.ceil is not None:
reward = min(reward, epoch.ceil)
if epoch.floor is not None:
reward = max(reward, epoch.floor)
epoch.add_reward(reward)
def adjust_hashrate(self, hashrate, epoch_idx, e_current, P_current, efficiency_current, electricity_cost_current):
alpha1 = None
alpha2 = None
alpha3 = None
alpha4 = None
# Found through regression on past values
epoch = self.epochs[epoch_idx]
log_eP_current = np.log(1+ (e_current * P_current))
log_eff_current = np.log(1+efficiency_current)
log_electricity_cost_current = np.log(1+electricity_cost_current)
# Put it all together:
[…]
predicted_log_hashrate = (
alpha1
+ alpha2 * log_eP_current
+ alpha3 * log_eff_current
+ alpha4 * log_electricity_cost_current
[…]
new_hashrate = np.exp(predicted_log_hashrate)
epoch.add_hashrate(new_hashrate)
return new_hashrate
def end_of_epoch(self):
last_epoch = self.epochs[-1]
P_B_median = last_epoch.median_block_reward()
N_n = last_epoch.average_hashrate()
new_epoch = Epoch()
BLOCKS_PER_EPOCH = 14
T_STAR = BLOCKS_PER_EPOCH * 10 * 60 # 1,209,600 for 2016 blocks
D_n = last_epoch.difficulty
T_n = (BLOCKS_PER_EPOCH * D_n * 2**32) / N_n
last_epoch.elapsed_time = T_n
T_hat = max(T_STAR /4, min(T_n, 4 *T_STAR))
D_next = D_n * (T_STAR / T_hat)
new_epoch.difficulty = D_next
# Case 1: Hashrate too high - need to decrease rewards
if self.DT > self.DT_N_UB:
#("Case 1: Hashrate above upper bound")
new_epoch.ceil = (self.tau) * P_B_median
new_epoch.floor = None # Clear any existing floor
# Case 2: Hashrate within bounds - gradual adjustment
elif self.DT_N_LB < self.DT < self.DT_N_UB:
# Previous epoch had a ceiling
if last_epoch.ceil is not None:
if last_epoch.ceil >= self.DT:
# Gradually relax the ceiling
new_epoch.ceil = (1+self.gamma) * last_epoch.ceil
else:
# Remove ceiling if it’s no longer needed
new_epoch.ceil = None
# Previous epoch had a floor
elif last_epoch.floor is not None:
if last_epoch.floor <= self.DT:
# Gradually relax the floor
new_epoch.ceil = (self.gamma) * last_epoch.floor
else:
# Remove floor if it’s no longer needed
new_epoch.floor = None
else:
new_epoch.floor = last_epoch.floor
new_epoch.ceil = last_epoch.ceil
# Case 3: Hashrate too low - need to increase rewards
elif self.DT < self.DT_N_LB:
#("Case 3: Hashrate below lower bound")
new_epoch.floor = (1+(1-self.tau)) * P_B_median
new_epoch.ceil = None # Clear any existing ceiling
self.epochs.append(new_epoch)
Appendix B Code for the Targeted Monetary Policy
This appendix states Python code for the Targeted monetary policy. The interpretation of the variables are commented. The code states the logic for the re-allocation of spending potential of addresses at each block that is induced by transactions and the addition or subtraction of the block reward, $p_{B_{n}}^{m}-p_{B_{n}}$, in accordance with the Hashrate Control Algorithm. It is not the code that would be written to implement the Targeted on servers in the Bitcoin network.
 2: Python code for Targeted monetary policy
def update_utxo_values(p_B_n_m, p_B_n, S_x_B_n_minus_1, P, FP, e_values, r_values, R, FR):
q = p_B_n_m - p_B_n
for x in range(len(e_values)):
if q < 0:
P += q
e_values[x] = math.floor(e_values[x] + S_x_B_n_minus_1[x] * (q + FP))
elif q > 0:
if P > q:
P -= q
e_values[x] = math.floor(e_values[x] - S_x_B_n_minus_1[x] * (q - FP))
else:
R += q - P
r_values[x] = math.floor(r_values[x] + S_x_B_n_minus_1[x] * (P + FR))
P = 0
e_values[x] = 0
x_values = [x_initial[i] + e_values[i] - r_values[i] for i in range(len(e_values))]
return P, e_values, r_values, R, FP, FR, x_values
def process_transactions(x_initial, e_values, r_values, transaction_indices):
# initial spending potential of addresses
x_values = [x_initial[i] + e_values[i] - r_values[i] for i in range(len(e_values))]
#transaction sender update step
for i in transaction_indices:
x_value = math.floor(x_initial[i] + e_values[i] - r_values[i])
e_values[i] = 0
r_values[i] = 0
#miner update step
x_value[miner_index] += p_B_n_m
total_sum = sum(x_values[i] + e_values[i] - r_values[i] for i in range(len(e_values)))
# Spending potential shares of addresses after transactions and miner fee, before allocations to UTXO Pool and Remainder Accounts
S_x_B_n_minus_1 = [(x_values[i] + e_values[i] - r_values[i]) / total_sum for i in range(len(e_values))]
return e_values, r_values, S_x_B_n_minus_1
def update_fp_fr(FP, FR, e_values, r_values):
FP += sum(e_values)
FR += sum(r_values)
return FP, FR
def compute_S_x_B_n(x_values, e_values, r_values):
total_sum = sum(x + e - r for x, e, r in zip(x_values, e_values, r_values))
# Spending potential shares of addresses after transactions and miner fee and assingments to UTXO Pool and Remainder Accounts.
S_x_B_n = [(x + e - r) / total_sum for x, e, r in zip(x_values, e_values, r_values)]
return S_x_B_n
Interpretation of variables:
x_initial = [<values to be inserted>] # initial values for x
e_initial = [<values to be inserted>] # initial UTXO values of addresses
r_initial = [<values to be inserted>] # initial values of r_x
p_B_n_m = <value to be inserted> # miner’s block reward
p_B_n = <value to be inserted> # total block reward
P = <value to be inserted> # aggregate value of UTXO Pool
e_values = e_initial[:] # copy of initial UTXO values in UTXO Pool
r_values = r_initial[:] # copy of initial values of r_x, the value for each address in the Remainder Accounts
R = <value to be inserted> # aggregate value in the Remainder Accounts
FP = <value to be inserted> # sum of values e_x in the UTXO Pool
FR = <values to be inserted> # sum of values r_x in the Remainder Accounts
transaction_indices = [:] # indices of addresses that send transactions in block B_n
miner_index = x # address of the miner of block B_N
Computation:
# Process transactions for each address in transaction_indices
e_values, r_values, S_x_B_n_minus_1 = process_transactions(x_initial, e_values, r_values, transaction_indices, p_B_n_m, miner_index)
# Update UTXO values for each address
P, e_values, r_values, R, FP, FR, x_values = update_utxo_values(p_B_n_m, p_B_n, S_x_B_n_minus_1, P, FP, e_values, r_values, R, FR)
# Update FP and FR with the residuals
FP, FR = update_fp_fr(FP, FR, e_values, r_values)
# Compute S_x_B_n for each address
S_x_B_n = compute_S_x_B_n(x_values, e_values, r_values)
# Compute the sum of initial and final x values
initial_x_sum = sum(x_initial)
final_x_sum = sum(x_values)
x_difference = initial_x_sum - final_x_sum
print("Updatedvalues:")
print(f"P:{P},e_values:{e_values},r_values:{r_values},x_values:{x_values},R:{R},FP:{FP},FR:{FR}")
print(f"S_x_B_n:{S_x_B_n}")
print(f"Differencebetweenthesumofinitialandfinalxvalues:{x_difference}")